OpenStudy (anonymous):

what is 3(k-2)=2(k+2)+k

OpenStudy (anonymous):

3(k-2) = 3k -6 2(k+2) = 2k +4 3k - 6 = 3k +4 -6 = 4 No solution

OpenStudy (anonymous):

u did it wrong

OpenStudy (anonymous):

3k - 6 = 2k + 4 >> here he wrote 3k + 4 by mistake ! so 3k - 2k = 4 + 6 k = 10 !!

OpenStudy (anonymous):

omg thanks

OpenStudy (anonymous):

thanks for the answer

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

ok i think she got the answer already!!:)

OpenStudy (anonymous):

daniel was correct. 2(k+2)+k=3k+4, there's no solution to the problem star posted.

OpenStudy (anonymous):

NOOOOOO there is an answer

OpenStudy (anonymous):

you agree that 2(k+2)+k=3k+4 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hey rsvitale there is an answer it was 10 just like atooty-s said AND IT IS RIGHT

OpenStudy (anonymous):

hello vernice thanks for giving me that problem it was awesome!!!!!

OpenStudy (anonymous):

AND DANIEL WAS WRONG

OpenStudy (anonymous):

then you typed the problem wrong. plug 10 in to what you stated. you get 3*8= 2*12 +10... 24=34. its wrong

OpenStudy (anonymous):

did you type the problem wrong

OpenStudy (anonymous):

she did

OpenStudy (anonymous):

oh ok so the +k isn't supposed to be there?

OpenStudy (anonymous):

never mind she did write it right

OpenStudy (anonymous):

I would know cuz we are in the same class right now.

OpenStudy (anonymous):

3(10-2)=2(10+2)+10 24=24 I DID RIGHT IT RIGHT AND I JUST ASKED MY TEACHER AND HE SAID 10 IS RIGHT

OpenStudy (anonymous):

LIAR

OpenStudy (anonymous):

calculate what you just typed. you get 24=34 youre doing the arithmetic wrong.

OpenStudy (anonymous):

BUT IT IS RIGHT OK

OpenStudy (anonymous):

DUDE ITS RIGHT!!!!! SHUTUP ALREADY

OpenStudy (anonymous):

STAR_PINK IS SOOO RUDE

OpenStudy (anonymous):

3*(10-2)=3*8=24... 2(10+2)+10=@*12+10=24+10=34. it is NOT right.

OpenStudy (anonymous):

OK WHATEVEER

OpenStudy (anonymous):

sorry 2*(10+2)=2*12=24..24+10=34

OpenStudy (anonymous):

WELL 34 AND 34 NOT 24 AND 34 LIKE YOU SAID

OpenStudy (anonymous):

3*8=24 not 34

OpenStudy (anonymous):

MY TEACHER JUST WORKED IT OUT !!

OpenStudy (anonymous):

if the problem is as you stated your teacher is wrong

OpenStudy (anonymous):

WELL YOUR TEACHER IS WRONG (RSVISTALE)

OpenStudy (anonymous):

WAT IS YOUR ANSWER :s

OpenStudy (anonymous):

there isn't a solution in the real numbers to that problem. If the problem is actually 3(k-2)=2(k+2) then k=10 is right.

OpenStudy (anonymous):

but the problem you posted doesn't have a solution

OpenStudy (anonymous):

OOOHHHHHHWWW BUT YOUR STILL WRONG

OpenStudy (anonymous):

why do you think i'm wrong

OpenStudy (anonymous):

BECAUSE YOU JUST ARE

OpenStudy (anonymous):

no, i'm not.

OpenStudy (anonymous):

YES YOU ARE

OpenStudy (anonymous):

OK THERE IS NO SOLUTION

OpenStudy (anonymous):

HOW OLD ARE YOU AND WHAT GRADE ARE YOU IN?

OpenStudy (anonymous):

im a phd student in math

OpenStudy (anonymous):

im 23

OpenStudy (anonymous):

WHAT IS THAT?

OpenStudy (anonymous):

doctoral degree. it's after college you continue study so that you can be a professor in university

OpenStudy (anonymous):

COOLL

OpenStudy (anonymous):

what grade are you guys in ?

OpenStudy (anonymous):

IM A GIRL IN 8TH GRADE

OpenStudy (anonymous):

HAHA

OpenStudy (anonymous):

the problem actually has a solution in some more advanced math but that's not anything you're talking about in 8th grade

OpenStudy (anonymous):

YES WE ARE WE ARE IN ADVANCE PRE-AP MATH THANK YOU

OpenStudy (anonymous):

THAT MEANS LIKE 10TH GRADE MATH

OpenStudy (anonymous):

i know i mean you cant solve something like that and get an answer until 3rd year of college.

OpenStudy (anonymous):

its called modular math

OpenStudy (anonymous):

anyway, cya later

OpenStudy (anonymous):

I DON'T CARE IF I WANT TO FIND A SOLUTION I CAN BY ASKING MY MOM SHE IS A TEACHER IN COLLAGE 4TH YEAR

OpenStudy (anonymous):

BYE GUTS I HAD SOO MUCH FUUN

OpenStudy (anonymous):

PEACE SAMANTHA SEE YOU AT LUNKCH